# ARS 548_ros: An ARS 548 RDI radar driver for ROS

### v3

@@ Abstract
The ARS 548 RDI Radar is a premium model of the fifth generation of 77 GHz long-range radar sensors with new RF antenna arrays, which offer digital beam forming. This radar measures independently the distance, speed, and angle of objects without any reflectors in one measurement cycle based on Pulse Compression with New Frequency Modulation [1]. Unfortunately, to the best of our knowledge, there are no open-source drivers available for Linux systems to enable users to analyze the data acquired by the sensor. In this paper, we present a driver that can interpret the data from the ARS 548 RDI sensor and make it available over the Robot Operating System versions 1 and 2 (ROS and ROS2). Thus, these data can be stored, represented, and analyzed using the powerful tools offered by ROS. Besides, our driver offers advanced object features provided by the sensor, such as relative estimated velocity and acceleration of each object, its orientation and angular velocity. We focus on the configuration of the sensor and the use of our driver including its filtering and representation tools. Besides, we offer a video tutorial to help in its configuration process. Finally, a dataset acquired with this sensor and an Ouster OS1-32 LiDAR sensor, to have baseline measurements, is available, so that the user can check the correctness of our driver [2].
keywords: ROS ; Linux driver ; Radar Sensor
@@
Legal Software License List one of the approved licenses
@@ 2 Software description
We developed our driver with two main objectives in mind. On the one hand, our driver generates standard ROS messages such as PointCloud2 that include information on the position and relative velocity, if available, of the detected points and objects; and PoseArray that indicates the direction of the movement of the detected objects. By using standard messages, the user can take advantage of existing representation and processing tools available in ROS. On the other hand, we are interested in providing the user with all the information generated by the radar without any processing. To this end, we have created the following custom ROS messages: ObjectList, DetectionList and Status (see Figure 2a), which are a direct translation of the data structures sent by the sensor as described in the ARS 548 reference manual [15]. The messages have been included in A for the sake of completeness. Finally, we can use the advanced information available in the Objects structure to filter them, so that only the required objects are taken into consideration (see Figure 2b and Section 2.2.3).
@@ 2.1 Software architecture
  • ARS 548_driver. This folder contains the libraries, code and functions needed for the actual driver to work. In particular, the ars_548 driver node receives the data from the radar via Ethernet link and transforms it into standard messages and custom messages (defined in ARS 548_messages), which are provided to the user via ROS middleware. On the other hand, ars_548_filter enables the user to distinguish different types of objects thanks to the advanced functionalities of the sensor. Finally, the radar_setup utility allows the user to change some parameters of the radar, such as its maximum detection range, its IP address, and many more.
@@
The full driver source code is publicly available in our GitHub repository [16]. Please refer to B for the additional contents associated with the paper.
@@ 2.2.3 Object filtering
The ARS 548_ros package provides the user with a filter able to interpret our custom messages and to provide filtered point clouds depending on the extended information generated by the sensor. The available information is defined in the Object message described in A. The filter can easily be customized via inheritance just by overriding a method; please refer to our repository for more details. As an example, see Section 3.3 in which we used a filter to represent only moving vehicles in the environment using the speed estimation from the sensor.
@@ 2.2.5 Radar configuration
We have included a basic utility that allows the user to change the configuration parameters of the ARS548. The configuration parameters of the device can be divided into the following categories. The reader is referred to the SensorConfiguration.msg file in A to see the detailed list of the configuration parameters:
@@
    Radar configuration. They change the mode of operation of the radar. You can configure its maximum detection distance (from 99 to 1500 meters), choose between three different frequency slots. They can also refer to internal configuration of the Radar, such as its IP address and power saving options, to name a few.
@@ 3.2 Driver Performance
We have performed a thirty-minute long experiment, in which we used our driver visualized the data from the ARS 548 sensor in real time with RViz. Our goal is to check the usual load and memory usage using the Ubuntu built-in system monitor. During the experiment, the driver process used a mere 2% of the processing power of one core. Furthermore, its memory usage was stable at around 7 MB. Therefore, the driver has a very low impact on the system resources.
@@ Appendix A Custom ARS 548 messages
Below you can find the definition of the custom ROS messages in ARS 548_messages.

### v4

@@ Abstract
The ARS 548 RDI Radar is a premium model of the fifth generation of 77 GHz long-range radar sensors with new RF antenna arrays, which offer digital beamforming. This radar measures independently the distance, speed, and angle of objects without any reflectors in one measurement cycle based on Pulse Compression with New Frequency Modulation. Unfortunately, to the best of our knowledge, there are no open-source drivers available for Linux systems to enable users to analyze the data acquired by the sensor. In this paper, we present a driver that can interpret the data from the ARS 548 RDI sensor and make it available over the Robot Operating System versions 1 and 2 (ROS and ROS2). Thus, these data can be stored, represented, and analyzed using the powerful tools offered by ROS. Besides, our driver offers advanced object features provided by the sensor, such as relative estimated velocity and acceleration of each object, its orientation and angular velocity. We focus on the configuration of the sensor and the use of our driver including its filtering and representation tools. Besides, we offer a video tutorial to help in its configuration process. Finally, a dataset acquired with this sensor and an Ouster OS1-32 LiDAR sensor, to have baseline measurements, is available so that the user can check the correctness of our driver.
keywords: ROS ; Linux driver ; Radar sensor ; ARS 548
@@
Legal Software License BSD 3-Clause License
@@ 2 Software description
We developed our driver with two main objectives in mind. On the one hand, our driver generates standard ROS messages such as PointCloud2 that include information on the position and relative velocity, if available, of the detected points and objects; and PoseArray that indicates the direction of the movement of the detected objects. By using standard messages, the user can take advantage of existing representation and processing tools available in ROS. On the other hand, we are interested in providing the user with all the information generated by the radar without any processing. To this end, we have created the following custom ROS messages: ObjectList, DetectionList and Status (see Figure 2a), which are a direct translation of the data structures sent by the sensor as described in the ARS 548 reference manual [15]. The messages have been included in Appendix A for the sake of completeness. Moreover, we can use the advanced information available in the Objects structure to filter them, so that only the required objects are taken into consideration (see Figure 2b and Section 2.2.3). Finally, we include the definition of the SensorConfiguration message, which is used in the configuration tool that can change the parameters of the ARS 548 RDI radar (see Section 2.2.5).
@@ 2.1 Software architecture
  • ars548_driver. This folder contains the libraries, code and functions needed for the actual driver to work. In particular, the ars548_driver driver node receives the data from the radar via Ethernet link and transforms it into standard messages and custom messages (defined in ars548_messages), which are provided to the user via ROS middleware. On the other hand, ars548_filter enables the user to distinguish different types of objects thanks to the advanced functionalities of the sensor. Finally, the radar_setup utility allows the user to change some parameters of the radar, such as its maximum detection range, its IP address, and many more.
@@
The full driver source code is publicly available in our GitHub repository [16]. Please refer to Appendix B for the additional contents associated with the paper.
@@ 2.2.3 Object filtering
The ars548_ros package provides the user with a filter able to interpret our custom messages and to provide filtered point clouds depending on the extended information generated by the sensor. The available information is defined in the Object message (see Appendix A). The filter can easily be customized via inheritance just by overriding a method; please refer to our repository for more details. As an example, see Section 3.3 in which we used a filter to represent only moving vehicles in the environment using the speed estimation from the sensor.
@@ 2.2.5 Radar configuration
We have included a basic utility that allows the user to change the configuration parameters of the ARS 548. The configuration parameters of the device can be divided into the following categories. The reader is referred to the SenorConfiguration.msg file in Appendix A to see the detailed list of the configuration parameters:
@@
    Radar configuration. They change the mode of operation of the radar. You can configure its maximum detection distance (from 99 to 1500 meters), and choose between three different frequency slots. They can also refer to internal configuration of the radar, such as its IP address and power saving options, to name a few.
@@ 3.2 Driver Performance
We have performed a thirty-minute long experiment, in which we used our driver and visualized the data from the ARS 548 sensor in real time with RViz. Our goal is to check the usual load and memory usage using the Ubuntu built-in system monitor. During the experiment, the driver process used a mere 2% of the processing power of one core. Furthermore, its memory usage was stable at around 7 MB. Therefore, the driver has a very low impact on the system resources.
@@ Appendix A Custom ARS 548 messages
Below you can find the definition of the custom ROS messages defined in the ars548_messages ROS package.